# Geometry-Aware Preference Learning for 3D Texture Generation

### v1

@@ Appendix D Additional Experiments and Results
We show expanded results from the main paper featuring more texture image and more viewpoints of different 3D objects. Figure 9, Figure 10, Figure 11, and Figure 12 shows the qualitative results of the symmetry-aware, geometry-texture alignment, geometry-guided texture colorization, and texture features emphasis experiments, respectively.
@@
Figure 9: Qualitative results of symmetry-aware experiment for different examples on a balloon mesh object. For each example (each rwo), we show the rendered 3D object from multiple viewpoints, alongside the corresponding texture images (rightmost column), which highlight the symmetric regions. A vertical dashed line marks the symmetry axis in each texture image. The purple plane passing through the center of the balloon in each viewpoint indicates the estimated symmetry plane of the object. As shown, compared to the pre-trained model (Tang et al., 2024), our method generates textures that are more consistent across symmetric parts of the mesh. Without symmetry supervision, patterns often differ noticeably between sides. In contrast, textures trained with the proposed symmetry reward exhibit visually coherent features across symmetric regions, demonstrating the reward’s effectiveness in enforcing symmetry consistency.
@@
Figure 10: Qualitative results of the geometry-texture alignment experiment on a rabbit (bunny) mesh. For each example, we show the rendered 3D object from multiple viewpoints, with the corresponding texture image in the rightmost column. Minimum curvature vectors, representing the underlying surface geometry, are visualized in the bottom row and overlaid on the textured objects for comparison. As shown, our method produces textures whose patterns align more closely with the mesh’s curvature directions, unlike InTex (Tang et al., 2024). Moreover, a notable outcome in our results is the emergence of repetitive texture patterns after fine-tuning with the geometry-texture alignment reward. This behavior arises from the differentiable sampling strategy used during reward computation. Specifically, it encourages the model to place edge features at specific UV coordinates which ultimately results in structured and repeated patterns in the texture (see Figure 5).
@@
Figure 12: Qualitative results of the texture features emphasis experiment on a rabbit (bunny) objects with different text prompts. For each example, we show the rendered 3D object from multiple viewpoints, with the corresponding texture image in the rightmost column. This goal is to learn texture images with salient features (e.g., edges) emphasized at regions of high surface bending, represented by the magnitude of mean curvature. This encourages texture patterns that highlight 3D surface structure while preserving perceptual richness through color variation. As illustrated, our method enhances texture features, such as edges and mortar, in proportion to local curvature, a capability InTex (Tang et al., 2024) lacks, often resulting in pattern-less (white) areas, particularly on the back and head of the rabbit.

### v2

@@ Appendix H Additional Experiments and Results
We show expanded results from the main paper featuring more texture image and more viewpoints of different 3D objects. Fig. 9, Fig. 10, Fig. 11, and Fig. 12 shows the qualitative results of the symmetry-aware, geometry-texture alignment, geometry-guided texture colorization, and texture features emphasis experiments, respectively. Moreover, a quantitative comparison on the bunny model is presented in Tab. 2, using various texture patterns. It evaluates our fine-tuned method against InTeX [30] (before fine-tuning) for geometry–texture alignment. Our approach yields significantly higher alignment counts, demonstrating that the generated textures respect the underlying geometric structure, whereas the InTeX [30] fails to achieve such alignment.
Table 2: A quantitative comparison on the bunny model, using various texture patterns, evaluates our fine‑tuned method against InTeX [30] (before fine‑tuning) for geometry–texture alignment. Each column reports the number of aligned vector pairs (i.e., minimum curvature directions versus texture gradients) in UV space. Our approach yields significantly higher alignment counts, demonstrating that the generated textures respect the underlying geometric structure, whereas the InTeX [30] baseline fails to achieve such alignment.
Texture pattern After fine-tune Before fine-tune
Tiger stripe 630/4487 (14.04%) 379/4487 (8.45%)
Black & white line 826/4487 (18.41%) 372/4487 (8.29%)
Checkerboard 485/4487 (10.81%) 321/4487 (7.15%)
Brick 477/4487 (10.63%) 302/4487 (6.73%)
Feather 556/4487 (12.39%) 343/4487 (7.64%)
Weave 472/4487 (10.52%) 299/4487 (6.66%)
Figure 8: Qualitative results of the camera-pose-aware texture generation experiment for different text prompts on different 3D mesh objects. The objective of this experiment is to learn optimal camera viewpoints such that, when the object is rendered and textured from these views, the resulting texture maximizes the average aesthetic reward. Consequently, by maximizing this reward, the model becomes invariant to the initial camera positions: regardless of where the cameras start, the training will adjust their azimuth and elevation to surround the 3D object in a way that yields high‑quality, aesthetically pleasing textures. We compare our fine‑tuning approach against InTeX [30] and TEXTure [26]. All methods share identical initial camera parameters; once initialized, each algorithm proceeds to paint the object. As shown, our method consistently outperforms both baselines, producing high‑fidelity, geometry‑aware textures across all shapes. This robustness stems from the aesthetic‑reward signal guiding camera adjustment during texture reward learning. In contrast, InTeX [30] and TEXTure [26] remain sensitive to poorly chosen initial viewpoints and thus often fail to generate coherent, high‑quality textures under suboptimal camera configurations.
@@
Figure 9: Qualitative results of symmetry-aware experiment for different examples on a balloon mesh object. For each example (each rwo), we show the rendered 3D object from multiple viewpoints, alongside the corresponding texture images (rightmost column), which highlight the symmetric regions. A vertical dashed line marks the symmetry axis in each texture image. The purple plane passing through the center of the balloon in each viewpoint indicates the estimated symmetry plane of the object. As shown, compared to the pre-trained model [30], our method generates textures that are more consistent across symmetric parts of the mesh. Without symmetry supervision, patterns often differ noticeably between sides. In contrast, textures trained with the proposed symmetry reward exhibit visually coherent features across symmetric regions, demonstrating the reward’s effectiveness in enforcing symmetry consistency.
@@
Figure 10: Qualitative results of the geometry-texture alignment experiment on a rabbit (bunny) mesh. For each example, we show the rendered 3D object from multiple viewpoints, with the corresponding texture image in the rightmost column. Minimum curvature vectors, representing the underlying surface geometry, are visualized in the bottom row and overlaid on the textured objects for comparison. As shown, our method produces textures whose patterns align more closely with the mesh’s curvature directions, unlike InTex [30]. Moreover, a notable outcome in our results is the emergence of repetitive texture patterns after fine-tuning with the geometry-texture alignment reward. This behavior arises from the differentiable sampling strategy used during reward computation. Specifically, it encourages the model to place edge features at specific UV coordinates which ultimately results in structured and repeated patterns in the texture (see Appendix D).
@@
Figure 12: Qualitative results of the texture features emphasis experiment on a rabbit (bunny) objects with different text prompts. For each example, we show the rendered 3D object from multiple viewpoints, with the corresponding texture image in the rightmost column. This goal is to learn texture images with salient features (e.g., edges) emphasized at regions of high surface bending, represented by the magnitude of mean curvature. This encourages texture patterns that highlight 3D surface structure while preserving perceptual richness through color variation. As illustrated, our method enhances texture features, such as edges and mortar, in proportion to local curvature, a capability InTex [30] lacks, often resulting in pattern-less (white) areas, particularly on the back and head of the rabbit.
Figure 13: Examples of four of the eleven questions in our user‐study questionnaire. We conducted a user‐preference study with 40 participants, each of whom completed 11 pairwise comparisons between textured 3D shapes generated by our method and by InTeX [30]. Participants rated visual quality according to four criteria: fidelity, symmetry, overall appearance, and text‑and‑texture pattern alignment. Specifically, for the three reward variants, cam‐pose‐aware texture generation, symmetry‐aware texture generation, and geometry–texture alignment, we asked three questions apiece, and for the texture‐features‐emphasis reward we asked two questions. We then applied one‐sided binomial tests to each experiment to obtain p‑values (see bottom of Tab. 1). As the table shows, the post‑reward‐learning results are significantly preferred over the pre‑learning outputs.